# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

### v3

@@ B.2 Dominated Integrability Lemma
B.3 Proofs of Self-Normalization Statements
@@
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.$ (121)
B.4 Proofs of Empirical Bernstein Statements
Proof of Proposition 4.10.
Recall that $\psi_{E}(\gamma)=-\log(1-\gamma)-\gamma$. An inequality by Fan et al. [2015] quoted by Howard et al. [2021, Appendix A.8] states that, for all $10\leqslant\gamma<$ 1 and $\xi\geqslant-1$,
$\exp(\gamma\xi-\psi_{E}(\gamma)\xi^{2})\leqslant 1+\gamma\xi.$ (122)
Since $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}\in\mathcal{S}_{d}^{[-1,\infty)}$, we can apply the transfer rule mentioned in Section B.3, replacing the scalar $\xi$ above by the matrix $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}$, and plugging in $\gamma=\gamma_{n}\in(0,1)$,
$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}).$ (123)
Lemma B.1 then guarantees the integrability of the left hand side, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbb{E}\left(1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\middle|\mathcal{F}_{n-1}\right)$ (124)
$=$ $1+\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (125)
where in the final step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. This concludes the proof. ∎
@@ B.4 Proofs of Empirical Bernstein Statements
We apply Lemma 5.1 with $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\gamma_{n}(\widehat{\mathbf{X}}_{n}-\mathbf{M}_{n})+\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})$ to see that
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)$ (126)
is a supermartingale, which upper bounds
$\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\right\}.$ (127)
Applying (UVI),
@@
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\geqslant\log(dU/\alpha)\right)\leqslant\alpha,$ (128)
and the desired result follows from rearranging. ∎
Lemma B.2.
Let $\psi_{E}(x)=-\log(1-x)-x$. When $0\leqslant x\leqslant\sqrt{2/5}$, $\psi_{E}(x)\leqslant x^{2}$.
Proof.
Let $g(x)=\psi_{E}(x)-x^{2}$. The claim follows from $g^{\prime\prime}(x)=(1-x)^{-2}-2\geqslant 0$ for $x\in[0,\sqrt{2/5}]$, and $g(0)=0$, $g(\sqrt{2/5})<0$. ∎
Proof of Corollary 5.7.
First, it is straightforward that $\lambda_{\max}({\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}})\geqslant-1$ for every $i=1,\dots,n$ since both $\mathbf{X}_{i}$ and $\overline{\mathbf{X}}_{i-1}$ take values in $\mathcal{S}_{d}^{[0,1]}$, so Theorem 5.6 is applicable. Let us prove the two claims about the deviation bound $D_{n}^{\mathsf{EB}}$ under $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$. Recall that
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (129)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right).$ (130)
First, via our definition, $\overline{v}_{i}\geqslant\frac{5\log(d/\alpha)}{n}$, therefore $\gamma_{i}\leqslant\sqrt{2/5}$. So Lemma B.2 implies that $\psi_{E}(\gamma_{i})\leqslant\gamma_{i}^{2}$. Therefore,
$D_{n}^{\mathsf{EB}}$ $=\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (131)
$\leqslant\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (132)
$=\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}.$ (133)
Then, let us show that $\lim_{n\to\infty}D_{n}^{\mathsf{EB}}/D_{n}^{\mathsf{B}}$ almost surely. Via the boundedness of $\mathbf{X}_{1}$ and the strong law of large numbers, we can see the following limits hold almost surely (cf. Waudby-Smith and Ramdas [2023, Lemmas 4-6]):
$\lim_{k\to\infty}\overline{\mathbf{X}}_{k}=\mathbf{M},\quad\lim_{k\to\infty}\overline{\mathbf{V}}_{k}=\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1}),$ (134)
$\lim_{k\to\infty}\overline{v}_{k}=\lambda_{\max}(\mathbf{V}),\quad\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}=\mathbf{V},$ (135)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}\right)=\lambda_{\max}(\mathbf{V}),$ (136)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)=1.$ (137)
Therefore, we have, via the expansion $\psi_{E}(x)=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$,
$\limsup_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}$ (138)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (139)
$\leqslant$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (140)
$+$ $\sum_{k=3}^{\infty}\limsup_{n\to\infty}\frac{\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{k}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (141)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (142)
$=$ $\limsup_{n\to\infty}\sqrt{\frac{\log(d/\alpha)}{2}}\frac{\left(1+\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)\right)}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}$ (143)
$=$ $\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}.$ (144)
Similarly, one can show that $\liminf_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}\geqslant\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$, concluding the proof. ∎
@@
Due to the monotonicity of $\log$ (Fact 2), the condition (68) implies
@@
Appendix C Further Notes on Matrix-Valued Martingales
C.1 Doob’s $L^{p}$ Inequality on Trace

### v4

@@ B.1 Dominated Integrability Lemma
B.2 Proofs for Forward Supermartingales
@@
Proof of Lemma 3.1.
Note that
$\{\mathbf{Y}_{n}\}$ is a supermartingale (99)
[…]
@@
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.\qed$ (112)
Proof of Example 3.12.
Using the transfer rule on the upper $p$-logarithmic contraction,
$\mathbb{E}\left(\exp(\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})))|\mathcal{F}_{n-1}\right)$ (113)
$\preceq$ $\mathbb{E}\left(\mathbf{I}+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+(1/p)\gamma_{n}^{p}\left(\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n})\right)^{p}|\mathcal{F}_{n-1}\right)$ (114)
$\preceq$ $\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n};$ (115)
and on the lower $p$-logarithmic contraction,
$\mathbb{E}\left(\exp(-\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})))|\mathcal{F}_{n-1}\right)$ (116)
$\preceq$ $\mathbb{E}\left(\mathbf{I}-\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+(1/p)\gamma_{n}^{p}\left(\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n})\right)^{p}|\mathcal{F}_{n-1}\right)$ (117)
$\preceq$ $\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}.$ (118)
This concludes the proof. ∎
B.3 Proofs for Backward Submartingales
Lemma B.2 (Contraction of the Matrix Variance).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$. Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (119)
Proof.
Note that
$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V}.\qed$ (120)
@@ B.3 Proofs for Backward Submartingales
Due to Fact 3,
Therefore,
is a scalar nonnegative backward submartingale. Using the scalar backward Ville’s inequality (7),
@@
$\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathbb{E}r_{1}=a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (123)
Finally, noting that $\|\mathbf{A}\|^{p}\leqslant\mathrm{tr}((\operatorname{abs}\mathbf{A})^{p})$ concludes the proof. ∎
B.4 Proof of Proposition 5.1
Recall for convenience the definition of sets
$\mathcal{C}_{\mathrm{tr}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathrm{tr}\mathbf{X}\leqslant d/\alpha\},$ (124)
$\mathcal{C}_{\mathbf{v}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant 1/\alpha\},$ (125)
$\mathcal{A}_{\alpha}=\{\mathbf{A}\in\mathcal{S}_{d}^{++}:\mathrm{tr}(\mathbf{A}^{-1})=\alpha\},$ (126)
$\mathcal{C}({\mathbf{A}})=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{X}\preceq\mathbf{A}\}.$ (127)
Proof of Statement 1.
Let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$. Then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}\mathbf{X}\leqslant d/\alpha$. Therefore $\mathbf{X}\in\mathcal{C}((d/\alpha)\mathbf{I})$. To see that there exists an $\mathbf{X}\in\mathcal{C}((d/\alpha)\mathbf{I})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can simply take $\mathbf{X}=(d/\alpha)\mathbf{I}$. ∎
Proof of Statement 2.
To see that there exists an $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can take $\mathbf{X}=\mathbf{A}$ as $\mathrm{tr}(\mathbf{A})\geqslant d^{2}/\alpha>d/\alpha$. To see that there exists an $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, suppose $\lambda_{\min}(\mathbf{A})=d/\alpha-2\delta$ where $\delta>0$. Take $\mathbf{X}$ to be a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$. ∎
Proof of Statement 3.
First, when $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}(\mathbf{X})\leqslant d/\alpha^{\prime}\leqslant 1/\lambda_{\max}(\mathbf{A}^{-1})=\lambda_{\min}(\mathbf{A})$ so $\mathbf{X}\in\mathcal{C}(\mathbf{A})$. Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$. ∎
Proof of Statement 4.
First, when $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}$, let $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ then $\mathrm{tr}\mathbf{X}\leqslant\mathrm{tr}\mathbf{A}\leqslant d/\alpha^{\prime}$ so $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$. Second, when $\alpha^{\prime}>d(\mathrm{tr}\mathbf{A})^{-1}$, take $\mathbf{X}=\mathbf{A}$ to see that $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ but $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$. ∎
Proof of Statement 5.
Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta=\frac{\alpha^{\prime}-\alpha}{d-1}$. ∎
Proof of Statement 6.
Let $\mathbf{A}$, $\alpha^{\prime}$, and $\mathbf{v}$ be arbitrary. Complete the orthonormal basis as in the previous proof. To find an $\mathbf{X}\in\mathcal{C}_{\mathbf{v}}(\alpha^{\prime})$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, take $\mathbf{X}={\alpha^{\prime}}^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}L\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for some sufficiently large $L$.
@@
Due to the monotonicity of $\log$ (Fact 1), the condition (67) implies
@@
Appendix C Miscellaneous Technical Results
C.1 Matrix MGF Bounds
In Table 1, we state several results in the random concentration literature on controlling the matrix MGF $\mathbf{G}(\gamma)=\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})})$ for a random matrix $\mathbf{X}_{1}$ satisfying certain conditions. If these conditions are met conditionally instead, bounds hold conditionally as well.
Name Condition $\mathbf{G}(\gamma)$
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
Bennett $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$
Bernstein $\mathbb{E}(\mathbf{X}_{1}-\mathbf{M})^{p}\preceq\frac{p!}{2}\mathbf{C}^{2}$ for $p=2,3,\dots$ $\preceq\mathrm{e}^{\frac{\gamma^{2}}{2(1-\gamma)}\mathbf{C}^{2}/2},\gamma\in(0,1)$
Table 1: Some important matrix MGFs. Here $\mathbf{C},\mathbf{V}$ are deterministic matrices, and $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$.
References for Table 1.
The uni-Gaussian MGF bound is proved by Tropp [2015, Lemma 4.6.2] via a Taylor expansion argument. For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section B.2). Bennett [Tropp, 2012, Lemma 6.7] and Bernstein [Tropp, 2012, Lemma 6.8] follow from a similar reasoning applied to the respective series expansion. ∎
C.2 A Matrix Doob’s $L^{p}$ Inequality